OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of p to the sixth power + the square root of 9 times p to the sixth power + 3 the square root of p to the sixth power − the square root of p squared 13 times p cubed times the square root of p − 2 times p times the square root of p 13 times p cubed times the square root of p − 2p 7p3 7p3 − p

OpenStudy (dumbcow):

ok easy way to think of these divide the power by the root, square root means the root is 2 6/2 = 3 \[\sqrt{p^{6}} = p^{6/2} = p^{3}\]